OpenStudy (meehan98):

In order to graph the following function: f(x)=log (subscript 2) (x+1), I know that from the parent function (log x), it's a horizontal shift to the left by 1, but I don't understand what to do with the subscript 2.

OpenStudy (freckles):

\[f(x)=\log_{2}(x+1) \\ f(x)=\frac{\log(x+1)}{\log(2)}=\frac{1}{\log(2)} \log(x+1)\] it just means you have to scale the parent function

OpenStudy (freckles):

1/log(2) is positive so it will keep the same direction if that makes sense to you

OpenStudy (mathmate):

recall: \(\large g(x)=a f(bx-h) + k\) a=vertical scale 1/b=horizontal scale h=horizontal (right) translation k=vertical (up) translation

OpenStudy (meehan98):

Ok, I think I understand that solution. In this equation: \[f(x)=\log _{3 }(\frac{ x }{ 4 })\] Do you follow the same steps? \[\frac{ \log \frac{ x }{ 4 } }{ \log3 }\]

OpenStudy (meehan98):

Never mind! I got it now!:)